# Reinforcement Learning and Machine ethics: a systematic review

### v1

@@ 2 Survey methodology
Objective function [74, 37]
Value function [2, 47, 63]
Others [15, 35, 51, 6]
@@
We can observe from these groupings that a majority of the articles modify either the reward function or the policy function. Reward functions act as means to instill ethical behavior, since the agent can obtain a higher reward or incur a lower penalty if it chooses the ethical option. On the other hand, inverse reinforcement learning algorithms are typically used to modify the policy function through demonstrations of an expert policy, based on which the agent learns the optimal reward function. Also, a few researchers demonstrated ethical behavior preference in RL agents by modifying either the value function or the neural network’s objective function (see Table 2). Other components include the action space [6], neural network [51] in case of deep RL algorithms, path planning [35] and regret [15].
RL paradigm
Similar to observing the component of the RL algorithm modified, it was also interesting to categorize and analyse the RL paradigm adopted to bring about ethical behavior. By paradigm, we refer to the type of RL used based on which the relevant component is modified. We briefly introduced some of the RL paradigms in Section Reinforcement Learning, such as inverse RL, multi-objective RL, constrained RL, etc. Each of these paradigms has variants and sub-variants, but these variations within paradigms largely function in a similar fashion. In Table 3, we specify the paradigms and the corresponding research works.
Reinforcement learning framework
@@
POMDP [65, 6]
Constrained/safe RL [21, 13, 18]
Regular RL [66, 3, 52, 50, 27, 12, 45, 25, 39, 76, 2, 37, 63, 15]
Other RL algorithms [60, 51, 74, 67, 23]
Table 3: RL frameworks used to result in ethical behavior.
@@
Most of the researchers adopted multi-objective RL (MORL) and inverse RL (IRL). The advantage of using MORL is that ethics specifications can be codified as part of the reward function, in addition to the main objective. An MORL agent navigates its environment, maximizing its reward thereby fulfilling its main tasks along with ethical sub-tasks, or in many cases, completes its main task in an ethical manner. While in IRL, the agent learns an optimal reward function based on an ethical policy demonstrated by an expert. Variations of IRL include, co-operative IRL, and adversarial IRL, which use human experts as a co-operative and adversarial means, respectively, rather than an expert teacher, to embed ethical behaviour in an artificial agent.
@@
Other techniques such as constrained RL and safe RL, teach an agent to ethically navigate an environment by avoiding unethical states or policies. Apart from these RL paradigms, several researchers also adopted regular RL, which refers to RL algorithms such as multi-arm bandits, deep Q-learning, actor-critic models, etc., which most applications of RL use. Finally, other RL algorithms include dynamic self-organizing maps [23], multi-agent RL [67], social RL [60], implicit RL [51] and affinity-based RL [74]. We also observe several uses of multi-agent RL, across these frameworks, however, for simplicity’s sake, we did not single out this categorization.
@@
A crucial aspect to explore in this systematic literature review is the ethical theories used by the researchers in their respective works. We reviewed each article to try to understand its underlying ethical proclivities. Based on this examination, we found that a majority does not explicitly state their ethics specifications and due to this, it was challenging to distinguish these manuscripts based on moral philosophy. However, to help researchers better understand the ethical underpinnings behind these machine ethics papers, we focus on the ethical theory that influences the RL algorithm the most, rather than all the theories at play. For example, if researchers codified social norms and then optimized a utility function as is done in RL, then we consider such implementations as deontological. On the other hand, if a paper does not explicitly mention deontological codification based on rules or norms, then it is considered consequentialist (which also includes variants of utilitarianism). We also found several other researchers using human values, where they optimized metrics based on human evaluation. In Table 4, we summarize the ethics distinctions.
Ethical theory
Consequentialist [55, 61, 67, 47, 57, 32, 63, 72, 32, 49, 25, 39]
Deontology [23, 31, 58, 45, 59, 37, 21, 6, 56, 38, 12, 18]
Virtue ethics [65, 74, 27, 53]
Human expert values [76, 23, 54, 69, 46, 24, 29, 36, 4, 64, 49, 13, 62]
Others (unspecified) [2, 3, 60, 51, 50]
@@
From this it can be observed that a majority adopted a consequentialist, deontological and human expert approach. Deontological and consequentialist ethics-based approaches being dominant is not surprising [65, 71], since they are arguably easier to codify and implement (compared to, for example, value ethics). We also see that there are several works under "human expert values", and that these "values" have been codified to help RL agents make ethical decisions in their environment. In an upcoming section, we deep-dive into who decides these values. Next, virtue ethics has fewer implementations, since the codification is challenging, however, it has been argued as a promising direction to pursue [74, 66]. Finally, the remaining manuscripts did not explicitly specify and instead had other goals such as social compliance [60], emulated imagination [50], etc.
Implementation Example
A wide variety of settings were used for experimental evaluation of systems. Very few of these settings were shared across teams of researchers revealing a lack of comparative benchmarking sets. As a result, ethical frameworks were generally compared against baselines in which there was no attempt to represent ethics, rather than alternative frameworks or theories. A broad categorisation of these examples can be seen in Table 5.
The only family of examples that appeared to have real traction in the literature were examples involving the equitable sharing of resources in a multi-agent setting – in which the metrics used were the ability of the community to support all its members. These scenarios ranged from abstract Grid Worlds representing a foraging games with various opportunities to share and/or steal resources from other agents, to more realistic settings based on, for instance, water distribution among several communities [35].
Another common family of examples were abstract representations of forbidden behaviour – often as a Grid World with forbidden areas [4] has a variant on this that involved the agent learning to play Tetris taking into account human preferences about colour placement. In many cases it was difficult to see these examples as specifically about ethical behaviour, as opposed to more general examples around safe reinforcement learning, learning preferences, or constrained learning.
Beyond these two distinct families of examples, there was a varied range of examples which involved motion planning in an environment that contained conflicting values or values conflicting with goals. For instance two papers [45, 46] required an agent to learn to play vegan pac-man where it was deemed unethical to eat the ghosts. Another set of papers all by the same team [55, 58, 57, 59] involved a “public civility game” in which agents need to reach a goal which is obstructed by rubbish and they must learn to divert to place the rubbish in a bin, rather than throwing it at the other agents involved.
Two examples involved moderately sophisticated simulations of smart grids [23] and financial markets [21].
The remaining cases were ad-hoc examples some with clear ethical relevance – e.g., around patient consent to treatment [38]. However, in some cases, particularly in systems where the focus was primarily on learning human norms, rather than a more specialised focus on values, we saw examples that seemed more about learning preferences than anything with particular ethical force – for instance [29] focuses on robot chef learning to cook meals that its owner will like.
Implementation example
Equity Example [66, 65, 67, 32, 56, 35, 27]
Abstract Example with Forbidden/Preferred States [37, 32, 51, 62, 12, 4]
Motion Planning with conflicting goals and/or values [76, 54, 45, 46, 55, 58, 57, 59, 49, 48]
Sophisticated Simulations [23, 21]
Other [31, 61, 2, 47, 15, 14, 38, 29, 25, 39]
Table 5: Categorisation of examples.
@@
We found several differences in claimed types of contributions in our literature review. Some were position papers, which were argumentative in nature, theorizing on how RL could be used in machine ethics research. Others were purely empirical contributions, either by referring to a previously proposed theory on a new problem, or improving upon existing results. The remaining ones were a combination of argumentative, empirical and mathematical proofs. It was interesting to explore these contribution types mainly to understand the maturity of the field: RL in machine ethics. The contributions are summarized in Table 6.
Contribution type
Argumentative/Theoretical [54, 6, 69, 52, 32, 36, 50, 27, 64, 74, 62, 18, 23, 66, 65, 67, 15, 24, 56, 51, 29, 48, 60, 72, 59]
Empirical [76, 61, 2, 47, 57, 37, 21, 49, 38, 35, 25, 39, 46, 4, 23, 66, 65, 67, 15, 24, 56, 51, 29, 48, 45, 63, 13, 59]
Mathematical Proof [60, 72, 45, 63, 13, 59]
@@
Usually, new fields begin with purely argumentative and position papers, discussing ways in which the research could move forward. Based on this, new mathematical and empirical evaluations are performed, and newer positions are formed. We can observe from the Table 6 that empirical evaluations form a large part of the research while still having a sizable number of theoretical contributions. In a later section, we discuss the research field as a whole and the associated trends based on contribution type.
Discussion
We here discuss the trends we observed in the surveyed papers and draw out some recommendations for future research based on our observations.
Recent increase of contributions in RL
Based on our survey, we found a steady increase in the use of RL in machine ethics. Figure 2 illustrates the trend, where we plot number of contributions per year. It is clear from this bar chart that since 2020, the attention has seen a consistent rise, with the peak in 2023.
Figure 2: Recent increase of contributions in RL based on the papers chosen for review.
No moral theories
Wallach [75] categorises approaches to the problem of machine ethics as either “top-down” or “bottom up”. Broadly, top down approaches seek to operationalise some ethical theory from Philosophy and apply this to a decision faced by the machine. Bottom-up approaches seek to learn ethical behaviour from data.
Much of the work covered in the survey papers [44, 71, 77] and performed in the early years of the field has taken a top down approach with a few systems taking a hybrid line – for instance the GenEth system [7] in which inputs from ethics experts were used to enable a system to learn explicit rules for differentiating between choices in a medical ethics situation. In reinforcement learning, the trend is not to follow a specific moral theory that will inform the agent’s policy. As we reflected in the previous section, there is no explicit intention to have reinforcement learning agent trained into following a specific moral theory.
Paradigm trends
A significant number of approaches we encountered have encoded the ethics in some fashion within the reward or objective function, within these we observed three trends:
@@ Discussion
In many cases the ethics signal was implemented as a reward signal in addition to the other (non-ethical) rewards such as maximising profit or efficiency. This frames ethical RL as a form of multi-objective RL.
@@
In other cases the ethics were represented as constraints. Ethical behaviour did not confer some reward upon the agent, although in some cases unethical behaviour incurred a penalty. This frames ethical RL as a form of constrained RL. In some cases there was an additional requirement that behaviour remain ethical during training in which case ethical RL becomes a form of Safe RL.
@@
Another significant theme, highlighted by the number of implementation examples based around equitable distribution of resources, is framing ethical ethical behaviour in terms of maximising group benefit in a multi-agent context.
@@
Human factors: whose ethics?
Based on our analysis of ethics specifications and moral philosophy, we found that several researchers have included a human factor to implement ethical behavior in RL agents. It is either the developer, the user, an expert or an adversary (Table 7), whose ethics an RL agent must try and emulate. Clearly, there is no real agreement on which ethical theory we, as humanity fall back on. We use some version of utilitarianism, social contracts, virtue ethics, deontology and other ethical theories interchangeably.
@@
Designer/developer/researcher [76, 23, 54, 14]
@@
Human expert/demonstrator [36, 46, 48]
@@
The problem of not clearly subscribing to a well understood moral theory is the human affinity for bias. Humans bias, is a hazard for designing applications that affect thousands of people who have their own plethora of biases that may be counter to the algorithm developer’s biases. Hence, it is crucial that algorithms consider a variety of viewpoints and evolve with changing times. While it is unrealistic to expect developers/product managers/researchers to know and understand so many viewpoints, it is still important that the makers strive towards at least an abstract understanding while developing ethical RL algorithms. Most of the algorithms in our survey have not explicitly specified who decides the ethics or right from wrong. However, we urge the research community to explicitly identify and report their ethical biases/assumptions made in their implementations, since future researchers who extend their work might be vulnerable to propagate the same.
@@
Some works (Table 7), where the human designer/developer/researcher decides ethical behavior, do indeed explicitly mention the human factor, where the RL algorithms are trained based on whether certain results are allowed or not. For instance, some of this is demonstrated using ethical movie recommendations, certain movies are excluded from the results for underage users [14]. While some use a judging RL agent trained with ethical values to train other RL agents in the environment [23]. An important challenge for such works is to codify human values and translate this to the algorithm. In cases where an expert human demonstrator is involved [36] in training the algorithm, it’s important that they are aware of competing ethical perspectives rather than training the model on one specific moral behavior. In the case of using co-operative humans and a human teammate, these are useful ways to train AI, but it’s also vital to infer what exactly is learnt by the model in the form of explanations and interpretations. Lastly, an interesting approach is to let the algorithm user specify the ethics [69]. This shifts the responsibility from the developer to the user, who can then customize the algorithm to their ethical needs. However, this approach could also affect the usability of the algorithm because with so much customization, the interaction experience might become troublesome for the user. Hence, striking the balance of usability and customization becomes important.
Summary

### v2

@@ 4.1 Component of RL Modified
Value function [39, 66, 67, 68]
Others [69, 70, 71, 72, 73, 74, 75, 76, 77, 22, 78]
@@
We can observe from these groupings that the majority of the articles modify either the reward function or the policy function. Reward functions act as a means to instill ethical behaviour, since the agent can obtain a higher reward or incur a lower penalty if it chooses the ethical option. On the other hand, inverse RL algorithms are typically used to modify the policy function through demonstrations of an expert policy, based on which the agent learns the optimal reward function. Also, a few researchers demonstrated ethical behaviour preference in RL agents by modifying either the value function or the neural network’s objective function (see Table 2). Other components include the action space [72], neural network [71] in the case of deep RL algorithms, path planning [70] and regret [69].
4.2 RL framework
Similar to observing the component of the RL algorithm modified, it was also interesting to categorise and analyse the RL framework adopted to bring about ethical behaviour. By framework, we refer to the type of RL used based on which the relevant component is modified. We briefly introduced some of the RL frameworks in Section 2.1, such as inverse RL, multi-objective RL, constrained RL, etc. Each of these frameworks has variants and sub-variants, but these variations within frameworks largely function in a similar fashion. In Table 3, we specify the frameworks and the corresponding research works.
RL framework
Regular RL [56, 40, 79, 80, 81, 82, 20, 83, 84, 43, 39, 78, 67, 69, 21, 58, 68, 85, 61, 76, 75, 74, 59, 73]
@@ 4.2 RL framework
Constrained/safe RL [48, 16, 89, 77]
Other RL algorithms [57, 71, 22, 42, 55, 72, 36, 60]
Table 3: RL frameworks used to result in ethical behaviour.
@@
Most of the researchers adopted multi-objective RL (MORL) and inverse RL (IRL). The advantage of using MORL is that ethics specifications can be codified as part of the reward function, in addition to the main objective. An MORL agent navigates its environment, maximising its reward, thereby fulfilling its main tasks along with ethical sub-tasks, or in many cases, completes its main task in an ethical manner. While in IRL, the agent learns an optimal reward function based on an ethical policy demonstrated by an expert. Variations of IRL include co-operative IRL, and adversarial IRL, which use human experts as co-operative and adversarial means, respectively, rather than an expert teacher, to embed ethical behaviour in an artificial agent.
@@
Other techniques, such as constrained RL and safe RL, teach an agent to ethically navigate an environment by avoiding unethical states or policies. Apart from these RL frameworks, several researchers have also adopted regular RL, which refers to RL algorithms such as multi-arm bandits, deep Q-learning, actor-critic models, etc., which most applications of RL use. Finally, other RL algorithms include dynamic self-organising maps [55], POMDP [36], multi-agent RL [42], social RL [57], implicit RL [71] and affinity-based RL [22]. We also observed several uses of multi-agent RL, across these frameworks; however, for simplicity’s sake, we did not single out this categorisation.
4.3 Implementation Example
A wide variety of settings were used for experimental evaluation of systems. Very few of these settings were shared across teams of researchers, revealing a lack of comparative benchmarking sets. As a result, ethical frameworks were generally compared against baselines in which there was no attempt to represent ethics, rather than alternative frameworks or theories. A broad categorisation of these examples can be seen in Table 4.
The only family of examples that appeared in various forms across multiple studies and teams of researchers involved the equitable sharing of resources in a multi-agent setting – in which the metrics used were the ability of the community to support all its members. These scenarios ranged from abstract Grid Worlds representing foraging games with various opportunities to share and/or steal resources from other agents, to more realistic settings based on, for instance, water distribution among several communities in Holgado-Sánchez et al. [70].
Another common family of examples was abstract representations of forbidden behaviour–often as a Grid World with forbidden areas. Alkoby et al. [44] has a variant on this that involved the agent learning to play Tetris taking into account human preferences about colour placement. In many cases, it was difficult to see these examples as specifically about ethical behaviour, as opposed to more general examples around safe RL, learning preferences, or constrained learning.
Beyond these two distinct families of examples, there was a varied range of examples which involved motion planning in an environment that contained conflicting values or values conflicting with goals. For instance, two papers,  Neufeld et al. [20] and Noothigattu et al. [50], required an agent to learn to play vegan pac-man where it was deemed unethical to eat the ghosts. Another set of papers, all by the same team Rodriguez-Soto et al. [46, 86, 87, 25] involved a “public civility game” in which agents need to reach a goal which is obstructed by rubbish, and they must learn to divert to place the rubbish in a bin, rather than throwing it at the other agents involved.
Two examples involved moderately sophisticated simulations of smart grids (Chaput et al. [55]) and financial markets (Byrd [48]).
The remaining cases were ad-hoc examples, some with clear ethical relevance – e.g., around patient consent to treatment (Kasenberg et al. [35]). However, in some cases, particularly in systems where the focus was primarily on learning human norms, rather than a more specialised focus on values, we saw examples that seemed more about learning preferences than anything with particular ethical force – for instance Fisac et al. [65] focuses on robot chef learning to cook meals that its owner will like.
Implementation example
Equity Example [56, 85, 42, 54, 19, 70, 81, 75]
Abstract Example with Forbidden/Preferred States [78, 54, 71, 47, 82, 44, 73, 77]
Motion Planning with conflicting goals and/or values [43, 62, 20, 50, 46, 86, 87, 25, 88, 41, 76, 74, 59]
Sophisticated Simulations [55, 48]
Other [53, 52, 39, 66, 69, 90, 35, 65, 83, 84, 36, 61, 60]
Table 4: Categorisation of examples.
4.4 Ethical Theory
A crucial aspect to explore in this systematic literature review is the ethical theories used by the researchers in their respective works. We reviewed each article to try to understand its underlying ethical proclivities. Based on this examination, we found that the majority do not explicitly state their ethics specifications and, due to this, it was challenging to distinguish these manuscripts based on moral philosophy. However, to help researchers better understand the ethical underpinnings behind these machine ethics papers, we focus on the ethical theory that influences the RL algorithm the most, rather than all the theories at play. For example, if researchers codified social norms and then optimised a utility function as is done in RL, then we consider such implementations as deontological. On the other hand, if a paper does not explicitly mention deontological codification based on rules or norms, then it is considered consequentialist (which also includes variants of utilitarianism). We also found several other researchers using human values, where they optimised metrics based on human evaluation. In Table 5, we summarise the ethics distinctions.
@@ 4.4 Ethical Theory
Consequentialist [46, 52, 42, 66, 87, 54, 67, 45, 54, 88, 83, 84, 21, 36, 59, 75]
Deontology [55, 53, 86, 20, 25, 78, 48, 72, 19, 35, 82, 89, 21, 77, 76, 74, 73]
Virtue ethics [85, 22, 81, 91, 21]
Human expert values [43, 55, 62, 51, 50, 49, 65, 64, 44, 63, 88, 16, 47, 58]
Others (unspecified) [39, 40, 57, 71, 80, 68, 60]
@@
From this it can be observed that a majority adopted a consequentialist, deontological and human expert approach. Deontological and consequentialist ethics-based approaches being dominant is not surprising [85, 11], since they are arguably easier to codify and implement (compared to, for example, virtue ethics). We also see that there are several works under ”human expert values”, and that these ”values” have been codified to help RL agents make ethical decisions in their environment. In an upcoming section, we deep-dive into who decides these values. Next, virtue ethics has fewer implementations, since the codification is challenging. However, it has been argued as a promising direction to pursue [22, 56]. Finally, the remaining manuscripts did not explicitly specify and instead had other goals such as social compliance [57], emulated imagination [80], etc.
4.5 Contribution type
We found several differences in claimed types of contributions in our literature review. Some were position papers, which were argumentative in nature, theorizing on how RL could be used in machine ethics research. Others were purely empirical contributions, either by referring to a previously proposed theory on a new problem, or improving upon existing results. The remaining ones were a combination of argumentative, empirical and mathematical proofs. It was interesting to explore these contribution types mainly to understand the maturity of the field: RL in machine ethics. The contributions are summarised in Table 6.
@@ 4.5 Contribution type
Argumentative/Theoretical [62, 72, 51, 79, 54, 64, 80, 81, 63, 22, 47, 89, 55, 56, 85, 42, 69, 49, 19, 71, 65, 41, 57, 45, 25, 68, 77, 60, 76, 75, 73, 36]
Empirical [43, 52, 39, 66, 87, 78, 48, 88, 35, 70, 83, 84, 50, 44, 55, 56, 85, 42, 69, 49, 19, 71, 65, 41, 20, 67, 16, 25, 21, 58, 61, 60, 76, 75, 74, 59, 36]
Math. Proof/Formalisation [57, 45, 20, 67, 16, 25, 77, 73, 36]
@@
Usually, new fields begin with purely argumentative and position papers, discussing ways in which the research could move forward. Based on this, new mathematical and empirical evaluations are performed, and newer positions are formed. We can observe from the Table 6 that empirical evaluations form a large part of the research while still having a sizeable number of theoretical contributions. In a later section, we discuss the research field as a whole and the associated trends based on contribution type.
5 Discussion
We here discuss the trends we observed in the surveyed papers and draw out some recommendations for future research based on our observations. We conclude this section by analysing research question RQ2 related to the human-factors responsible for deciding the ethics.
5.1 Recent increase in
contributions in RL Based on our survey, we found a steady increase in the use of RL in machine ethics. Figure 2 illustrates the trend, where we plot the number of contributions per year. It is clear from this bar chart that, since 2020, the attention towards ME and RL has seen a consistent rise, with a peak in 2022.
Figure 2: Recent increase in contributions in RL based on the papers chosen for review.
5.2 No moral theories
Wallach [3] categorises approaches to the problem of machine ethics as either “top-down” or “bottom up”. Broadly, top-down approaches seek to operationalise some ethical theory from philosophy and apply this to a decision faced by the machine. Bottom-up approaches seek to learn ethical behaviour from data.
Much of the work covered in the survey papers [10, 11, 12] and performed in the early years of the field has taken a top-down approach with a few systems taking a hybrid line – for instance, the GenEth system [92] in which inputs from ethics experts were used to enable a system to learn explicit rules for differentiating between choices in a medical ethics situation. In RL, the trend is not to follow a specific moral theory that will inform the agent’s policy. As we reflected in the previous section, there is no explicit intention to have RL agents trained into following a specific moral theory.
5.3 Paradigm trends
A significant number of approaches we encountered have encoded the ethics in some fashion within the reward or objective function. Within these we observed three trends:
@@ 5.3 Paradigm trends
In many cases, the ethics was implemented as a reward signal in addition to other (non-ethical) rewards, such as maximising profit or efficiency. This frames ethical RL as a form of multi-objective RL.
@@
In other cases, the ethics were represented as constraints. Ethical behaviour does not confer any reward upon the agent, although in some cases unethical behaviour incurred a penalty. This frames ethical RL as a form of constrained RL. In some cases, there is an additional requirement that behaviour remain ethical during training, in which case ethical RL becomes a form of Safe RL.
@@
Another significant theme, highlighted by the number of implementation examples based around equitable distribution of resources, is framing ethical behaviour in terms of maximising group benefit in a multi-agent context.
@@
5.4 Human factors: whose ethics?
Based on our analysis of ethics specifications and moral philosophy, we found that several researchers have included a human factor to implement ethical behaviour in RL agents. It is either the developer, the user, an expert, or an adversary (Table 7), whose ethics an RL agent must try and emulate. Clearly, there is no real agreement on which ethical theory we, as humanity, fall back on. We use some version of utilitarianism, social contracts, virtue ethics, deontology, and other ethical theories interchangeably.
@@ 5.4 Human factors: whose ethics?
Designer/developer/researcher [69, 90, 52, 78, 71, 82, 19, 42, 56, 81, 70, 39, 54, 83, 43, 85, 46, 86, 87, 25, 88, 35, 48, 55, 66, 62, 20, 40, 45, 80, 22, 89, 21, 61, 77, 76, 75, 74, 59, 73, 36]
@@
Human expert/demonstrator [64, 50, 41, 58]
@@
The problem of not clearly subscribing to a well-understood moral theory is the human affinity for bias. Human bias is a hazard for designing applications that affect thousands of people who have their own plethora of biases that may be counter to the algorithm developer’s biases. Hence, it is crucial that algorithms consider a variety of viewpoints and evolve with changing times. While it is unrealistic to expect developers/product managers/researchers to know and understand so many viewpoints, it is still important that the makers strive towards at least an abstract understanding while developing ethical RL algorithms. Most of the algorithms in our survey have not explicitly specified who decides the ethics or right from wrong. However, we urge the research community to explicitly identify and report their ethical biases/assumptions made in their implementations, since future researchers who extend their work might be vulnerable to propagating the same [93].
@@
Some works (Table 7), where the human designer/developer/researcher decides ethical behaviour, do indeed explicitly mention the human factor, where the RL algorithms are trained based on whether certain results are allowed or not. For instance, some of this is demonstrated using ethical movie recommendations, where certain movies are excluded from the results for underage users [90]. Others [55] use a judging RL agent trained with ethical values to train other RL agents in the environment. An important challenge for such works is to codify human values and translate these into the algorithm. In cases where an expert human demonstrator is involved [64] in training the algorithm, it’s important that they are aware of competing ethical perspectives rather than training the model on one specific moral behaviour. In the case of using co-operative humans and a human teammate, these are useful ways to train AI, but it’s also vital to infer what exactly is learnt by the model in the form of explanations and interpretations. Lastly, an interesting approach is to let the algorithm user specify the ethics [51]. This shifts the responsibility from the developer to the user, who can then customise the algorithm to their ethical needs. However, this approach could also affect the usability of the algorithm because, with so much customisation, the interaction experience might become troublesome for the user. Hence, striking the balance of usability and customisation becomes important.
6 Summary